A stone is thrown into the air with initial speed of 3m/s from the ground.


What time does it hit the ground? (Neglect air resistance)


1
Expert's answer
2022-03-23T07:15:51-0400

The time elapsed as a stone was thrown and it hit the ground

"\Delta t=2\times \frac{v_0}{g}\\
=\rm 2\times\frac{3.0\: m/s}{9.8\: m/s^2}=0.61\: s"